# Do Foundational Audio Encoders Understand Music Structure?

### v1

@@ 4.2 Discussions
| OpenL3 | ||||||||
@@ 5 Conclusion
Although many open-source foundational audio encoders (FAEs) exist, their capabilities in understanding music structure remain underexplored. In this study, we conducted a comparative evaluation of 11 types of FAEs, considering factors such as learning methods, training data, and model context length. Our experimental results demonstrate that self-supervised masked language models trained on music data achieve strong performance, and that a longer context length for FAEs yields better results. This finding aligns with the requirements of the music structure analysis task, which involves capturing long-term musical contexts. Our results contribute to future research on music structure analysis and FAEs. The strong FAEs identified in this work can serve as backbones for metrics such as Fréchet Audio Distance (FAD) to evaluate long-form music generation models.

### v2

@@ 4.1 Results
| OpenL3 | 38.331.24 | 22.650.86 | 50.240.95 | 44.481.20 | 60.301.88 | 60.151.05 | 58.092.40 | 58.451.23 |
4.2 Discussion
Pooling. MLM models confirmed the effectiveness of pooling. HR3F improved across all MLM models, and PWF improved with MERT, while the two AudioMAEs improved across all metrics. Nevertheless, pooling can cause boundary position shifts, indicating that HR.5F may degrade for some methods.
Feature frame rate. MULE, an SSL model using contrastive learning, showed low boundary detection performance, likely due to its coarse frame rate of 0.5 Hz and the resulting blurred boundaries, as shown in Fig. 1. However, an excessively high frame rate, such as in the case of EnCodec and DAC, might also harm boundary detection. Among the best three models, MERT (75 Hz) and MusicFM (25 Hz) have a high frame rate, but AudioMAE (Zhong) has only 6.25 Hz, indicating a high frame rate is not essential for strong MSA performance.
Semantic vs. acoustic SSL. The tokenizers (codecs) and MLM models are similar in their training objectives, as both require models to reconstruct the input signal. However, tokenizers performed very poorly in both boundary detection and function prediction. We believe the following differences lead to the observed performance gap. MLM encourages models to reconstruct original signals from a masked version without pursuing reconstruction details, resulting in semantic features. In contrast, tokenizers are trained to recover as many acoustic details as possible, rather than semantic ones. In Fig. 1, numerous fine vertical stripes are visible for EnCodec and DAC when pooling is not applied, which is the characteristic of fine-grained details. Pooling reduces these vertical stripes, but blurs boundaries and further degrades performance in these models. We also examined EnCodec with various bit rates, but observed no statistically meaningful performance difference.
Architecture. PANNs (CNN) and PaSST (Transformer), as well as OpenL3 (CNN) and CLAP (Transformer), which share similar learning methods, achieve similar performance, indicating that model architecture is not the key factor for these methods. However, for longer 30-second training data, as with MusicFM, an architecture with a long receptive field, such as Transformers, might be preferred.
Models used as metrics to evaluate music generation quality. The following FAEs have been used as backbones in metrics such as FAD to evaluate the quality of music and audio generative models: MERT in [12, 14]; EnCodec, DAC, and CLAP in [12]; PANNs in [22]; PaSST in [43]; and OpenL3, which has been used in long-form music evaluation [10]. In this study, we found that, compared to most of the aforementioned models, MLM models, particularly MusicFM, exhibit a better understanding of music structure. We encourage the community to revisit the validity of these metrics and hypothesize that MLM methods could serve as strong alternatives. In a recent metric called MAD [14], MERT has been chosen as the backbone, which we believe is a positive sign.
Future work. An important branch of SSL is auto-regression (AR), which is not included in this research, as we found that AR models, such as Jukebox [9] and SoniDo [21], are resource-intensive due to their more-than-5B-parameters. It would be interesting to explore AR models at a scale similar to MLM models (less than 330M parameters) and evaluate their potential in the MSA task. We also noticed that, instead of a linear probe, backends utilizing longer contexts have been shown to improve MSA performance [41]. Moreover, different layers in Transformer-based FAEs tend to contain different features [9]. Investigating features from different FAE layers and designing more powerful backends are promising directions for future work.
@@ 5 Conclusion
This study addressed the underexplored capabilities of foundational audio encoders (FAEs) for music structure analysis (MSA). Through a comparative evaluation of 11 FAEs, we investigated key factors such as learning method, training data, and context length. Our experimental results demonstrate that self-supervised masked language models trained on music data with a long context length achieve the strongest performance, a finding consistent with MSA’s inherent need to capture long-term musical context. Our findings provide clear guidance for future research on MSA and FAEs, suggesting that the strong FAEs identified in this study can also serve as improved backbones for metrics such as FAD to evaluate long-form music generation.